# The fractional $p$-Laplacian on hyperbolic spaces

### v2

@@ 4. Extension problem
$P(\rho,y)=C_{4}\,y^{sp}\left(\frac{-\partial_{\rho}}{\sinh\rho}\right)^{\frac{n-1}{2}}\mathscr{K}_{\frac{1+sp}{2},\frac{n-1}{2}}\left(\sqrt{\rho^{2}+y^{2}}\right)$
@@
$P(\rho,y)=C_{4}\,y^{sp}\int_{\rho}^{\infty}\frac{\sinh r}{\sqrt{\pi}\sqrt{\cosh r-\cosh\rho}}\left(\frac{-\partial_{r}}{\sinh r}\right)^{\frac{n}{2}}\mathscr{K}_{\frac{1+sp}{2},\frac{n-1}{2}}\left(\sqrt{r^{2}+y^{2}}\right)\,\mathrm{d}r$
when $n\geq 2$ even; here, $\mathscr{K}_{\nu,a}$ is the function given in (2.4) and
@@
$V(x,y)=\frac{y^{sp}}{2^{sp}\Gamma(\frac{sp}{2})}\int_{0}^{\infty}\int_{\mathbb{H}^{n}}p(t,\rho)f(\xi)\,\mathrm{d}\xi\,e^{-\frac{y^{2}}{4t}}\frac{\mathrm{d}t}{t^{1+\frac{sp}{2}}},$
where $\rho=d(x,\xi)$. Recalling the expression (3.1) for the heat kernel $p(t,\rho)$ and using (2.1), we obtain
@@
$\begin{split}\int_{0}^{\infty}e^{-\frac{(2m-1)^{2}}{4}t-\frac{r^{2}+y^{2}}{4t}}t^{-\frac{5+sp}{2}}\,\mathrm{d}t&=2(2m-1)^{\frac{3+sp}{2}}\mathscr{K}_{\frac{3+sp}{2},\frac{2m-1}{2}}\left(\sqrt{r^{2}+y^{2}}\right)\\ &=4(2m-1)^{\frac{1+sp}{2}}\left(\frac{-\partial_{r}}{r}\right)\mathscr{K}_{\frac{1+sp}{2},\frac{2m-1}{2}}\left(\sqrt{r^{2}+y^{2}}\right).\end{split}$
@@
$\begin{split}V(x,y)&=C_{4}\,y^{sp}\int_{\mathbb{H}^{n}}\left(\frac{-\partial_{\rho}}{\sinh\rho}\right)^{m-1}\int_{\rho}^{\infty}\frac{(-\partial_{r})\mathscr{K}_{\frac{1+sp}{2},\frac{2m-1}{2}}\left(\sqrt{r^{2}+y^{2}}\right)}{\sqrt{\pi}\sqrt{\cosh r-\cosh\rho}}\,\mathrm{d}r\,f(\xi)\,\mathrm{d}\xi\\ &=\int_{\mathbb{H}^{n}}P(d(x,\xi),y)f(\xi)\,\mathrm{d}\xi\end{split}$
@@
Finally, we prove $V(x,0)=f(x)$. Indeed, we have $P(\rho,y)\to 0$ as $y\searrow 0$ if $\rho\neq 0$ by definition. Moreover, since the heat kernel $p(t,\rho)$ satisfies
$\int_{\mathbb{H}^{n}}p(t,d(x,\xi))\,\mathrm{d}\xi=1,$
@@
$\begin{split}\int_{\mathbb{H}^{n}}P(d(x,\xi),y)\,\mathrm{d}\xi&=\frac{y^{sp}}{2^{sp}\Gamma(\frac{sp}{2})}\int_{0}^{\infty}\left(\int_{\mathbb{H}^{n}}p(t,d(x,\xi))\,\mathrm{d}\xi\right)e^{-\frac{y^{2}}{4t}}\frac{\mathrm{d}t}{t^{1+\frac{sp}{2}}}\\ &=\frac{y^{sp}}{2^{sp}\Gamma(\frac{sp}{2})}\int_{0}^{\infty}e^{-\frac{y^{2}}{4t}}\frac{\mathrm{d}t}{t^{1+\frac{sp}{2}}}=1.\end{split}$
This concludes that $V$ solves the extension problem (1.4). ∎
@@
$K(\rho)=\left(\frac{-\partial_{\rho}}{\sinh\rho}\right)^{\frac{n-1}{2}}\left(\mathscr{K}_{\frac{1+sp}{2},\frac{n-1}{2}}(\rho)-\mathscr{K}_{\frac{1+sp}{2},\frac{n-1}{2}}\left(\sqrt{\rho^{2}+y^{2}}\right)\right)$
@@
$K(\rho)=\int_{\rho}^{\infty}\frac{\sinh r}{\sqrt{\pi}\sqrt{\cosh r-\cosh\rho}}\left(\frac{-\partial_{r}}{\sinh r}\right)^{\frac{n}{2}}\left(\mathscr{K}_{\frac{1+sp}{2},\frac{n-1}{2}}(r)-\mathscr{K}_{\frac{1+sp}{2},\frac{n-1}{2}}\left(\sqrt{r^{2}+y^{2}}\right)\right)\mathrm{d}r$
@@ 5. Pointwise convergence
as $s\to 1^{-}$, play a key role in the proof of Theorem 1.5.
@@
Let us first consider the case $n=2m+1$ with $m\geq 1$. Since $c_{n,s,p}C_{1}\leq C(1-s)$ for some $C=C(n,p)>0$, by using Corollary 3.3 we have
(5.3) $\begin{split}0&\leq c_{n,s,p}\int_{R}^{\infty}\mathcal{K}_{n,s,p}(\rho)\sinh^{n-1}\rho\,\mathrm{d}\rho\\ &\lesssim(1-s)\int_{R}^{\infty}\sinh^{2m}\rho\left(\frac{-\partial_{\rho}}{\sinh\rho}\right)^{m}\mathscr{K}_{\frac{1+sp}{2},m}(\rho)\,\mathrm{d}\rho.\end{split}$
@@
(5.4) $\lim_{s\nearrow 1}(1-s)\int_{R}^{\infty}\sinh^{m+a}\rho\left(\frac{-\partial_{\rho}}{\sinh\rho}\right)^{m}\mathscr{K}_{\frac{1+sp}{2},a}(\rho)\,\mathrm{d}\rho=0\quad\text{for each }a>0.$
@@
$\begin{split}\int_{R}^{\infty}\sinh^{1+a}\rho\left(\frac{-\partial_{\rho}}{\sinh\rho}\right)\mathscr{K}_{\frac{1+sp}{2},a}(\rho)\,\mathrm{d}\rho&=a\int_{R}^{\infty}(\sinh^{a}\rho)\rho^{-\frac{1+sp}{2}}K_{\frac{3+sp}{2}}(a\rho)\,\mathrm{d}\rho\\ &\leq a\int_{R}^{\infty}(\sinh^{a}\rho)\rho^{-\frac{1+sp}{2}}K_{\frac{3+p}{2}}(a\rho)\,\mathrm{d}\rho.\end{split}$
@@
$\lim_{s\nearrow 1}(1-s)\int_{R}^{\infty}\sinh^{1+a}\rho\left(\frac{-\partial_{\rho}}{\sinh\rho}\right)\mathscr{K}_{\frac{1+sp}{2},a}(\rho)\,\mathrm{d}\rho=0,$
@@
$\int_{R}^{\infty}\sinh^{m+1+a}\rho\left(\frac{-\partial_{\rho}}{\sinh\rho}\right)^{m+1}\mathscr{K}_{\frac{1+sp}{2},a}(\rho)\,\mathrm{d}\rho$
$=(m+a)\int_{R}^{\infty}\sinh^{m+a-1}\rho\cosh\rho\left(\frac{-\partial_{\rho}}{\sinh\rho}\right)^{m}\mathscr{K}_{\frac{1+sp}{2},a}(\rho)\,\mathrm{d}\rho$
$\quad-\left[\sinh^{m+a}\rho\left(\frac{-\partial_{\rho}}{\sinh\rho}\right)^{m}\mathscr{K}_{\frac{1+sp}{2},a}(\rho)\right]_{R}^{\infty}.$
Thus, taking the limit $\lim_{s\nearrow 1}(1-s)$ and using Lemma 3.4 yields
$\begin{split}&\lim_{s\nearrow 1}(1-s)\int_{R}^{\infty}\sinh^{m+1+a}\rho\left(\frac{-\partial_{\rho}}{\sinh\rho}\right)^{m+1}\mathscr{K}_{\frac{1+sp}{2},a}(\rho)\,\mathrm{d}\rho\\ &=\lim_{s\nearrow 1}(1-s)(m+a)\int_{R}^{\infty}\sinh^{m+a-1}\rho\cosh\rho\left(\frac{-\partial_{\rho}}{\sinh\rho}\right)^{m}\mathscr{K}_{\frac{1+sp}{2},a}(\rho)\,\mathrm{d}\rho.\end{split}$
@@
$\begin{split}&\lim_{s\nearrow 1}(1-s)\int_{R}^{\infty}\sinh^{m+1+a}\rho\left(\frac{-\partial_{\rho}}{\sinh\rho}\right)^{m+1}\mathscr{K}_{\frac{1+sp}{2},a}(\rho)\,\mathrm{d}\rho\\ &\leq(m+a)(\coth R)\lim_{s\nearrow 1}(1-s)\int_{R}^{\infty}\sinh^{m+a}\rho\left(\frac{-\partial_{\rho}}{\sinh\rho}\right)^{m}\mathscr{K}_{\frac{1+sp}{2},a}(\rho)\,\mathrm{d}\rho=0.\end{split}$
@@
$\begin{split}0&\leq c_{n,s,p}\int_{R}^{\infty}\mathcal{K}_{n,s,p}(\rho)\sinh^{n-1}\rho\,\mathrm{d}\rho\\ &\lesssim(1-s)\int_{R}^{\infty}\sinh^{2m-1}\rho\int_{\rho}^{\infty}\frac{\sinh r}{\sqrt{\cosh r-\cosh\rho}}\left(\frac{-\partial_{r}}{\sinh r}\right)^{m}\mathscr{K}_{\frac{1+sp}{2},\frac{2m-1}{2}}(r)\,\mathrm{d}r\,\mathrm{d}\rho,\end{split}$
@@
(5.7) $\begin{split}\lim_{s\nearrow 1}(1-s)\int_{R}^{\infty}\sinh^{\frac{2m-1}{2}+a}\rho\int_{\rho}^{\infty}\frac{\sinh r}{\sqrt{\cosh r-\cosh\rho}}\left(\frac{-\partial_{r}}{\sinh r}\right)^{m}\mathscr{K}_{\frac{1+sp}{2},a}(r)\,\mathrm{d}r\,\mathrm{d}\rho=0\\ \quad\text{for each }a\geq 1/2.\end{split}$
@@
$\begin{split}&\int_{R}^{\infty}\sinh^{\frac{1}{2}+a}\rho\int_{\rho}^{\infty}\frac{\sinh r}{\sqrt{\cosh r-\cosh\rho}}\left(\frac{-\partial_{r}}{\sinh r}\right)\mathscr{K}_{\frac{1+sp}{2},a}(r)\,\mathrm{d}r\,\mathrm{d}\rho\\ &\leq a\int_{R}^{M/a}\sinh^{\frac{1}{2}+a}\rho\int_{\rho}^{\infty}\frac{1}{\sqrt{\cosh r-\cosh\rho}}r^{-\frac{1+sp}{2}}K_{\frac{3+p}{2}}(ar)\,\mathrm{d}r\,\mathrm{d}\rho\\ &\quad+a\int_{M/a}^{\infty}\sinh^{\frac{1}{2}+a}\rho\int_{\rho}^{\infty}\frac{1}{\sqrt{\cosh r-\cosh\rho}}r^{-\frac{1+sp}{2}}K_{\frac{3+p}{2}}(ar)\,\mathrm{d}r\,\mathrm{d}\rho=:J_{1}+J_{2}.\end{split}$
@@
$\begin{split}&\int_{R}^{\infty}\sinh^{\frac{2m+1}{2}+a}\rho\int_{\rho}^{\infty}\frac{\sinh r}{\sqrt{\cosh r-\cosh\rho}}\left(\frac{-\partial_{r}}{\sinh r}\right)^{m+1}\mathscr{K}_{\frac{1+sp}{2},a}(r)\,\mathrm{d}r\,\mathrm{d}\rho\\ &=\int_{R}^{\infty}\sinh^{\frac{2m+1}{2}+a}\rho\left(\frac{-\partial_{\rho}}{\sinh\rho}\right)\int_{\rho}^{\infty}\frac{\sinh r}{\sqrt{\cosh r-\cosh\rho}}\left(\frac{-\partial_{r}}{\sinh r}\right)^{m}\mathscr{K}_{\frac{1+sp}{2},a}(r)\,\mathrm{d}r\,\mathrm{d}\rho.\end{split}$
@@
$\begin{split}&\lim_{s\nearrow 1}(1-s)\int_{R}^{\infty}\sinh^{\frac{2m+1}{2}+a}\rho\int_{\rho}^{\infty}\frac{\sinh r}{\sqrt{\cosh r-\cosh\rho}}\left(\frac{-\partial_{r}}{\sinh r}\right)^{m+1}\mathscr{K}_{\frac{1+sp}{2},a}(r)\,\mathrm{d}r\,\mathrm{d}\rho\\ &\leq C\lim_{s\nearrow 1}(1-s)\int_{R}^{\infty}\sinh^{\frac{2m-1}{2}+a}\rho\int_{\rho}^{\infty}\frac{\sinh r}{\sqrt{\cosh r-\cosh\rho}}\left(\frac{-\partial_{r}}{\sinh r}\right)^{m}\mathscr{K}_{\frac{1+sp}{2},a}(r)\,\mathrm{d}r\,\mathrm{d}\rho\end{split}$
@@
(5.9) $\begin{split}&\lim_{s\nearrow 1}(1-s)\int_{0}^{R}\rho^{p}\sinh^{2m}\rho\left(\frac{-\partial_{\rho}}{\sinh\rho}\right)^{m}\mathscr{K}_{\frac{1+sp}{2},m}(\rho)\,\mathrm{d}\rho\\ &=\frac{2^{m-1}}{p}\left(\frac{2}{m}\right)^{\frac{p+1}{2}}\Gamma\left(\frac{p+2m+1}{2}\right)\end{split}$
@@
(5.10) $\begin{split}\lim_{s\nearrow 1}(1-s)\int_{0}^{R}\rho^{p}\sinh^{2m}\rho\left(\frac{-\partial_{\rho}}{\sinh\rho}\right)^{m}\mathscr{K}_{\frac{1+sp}{2},a}(\rho)\,\mathrm{d}\rho\\ =\frac{2^{m-1}}{p}\left(\frac{2}{a}\right)^{\frac{p+1}{2}}\Gamma\left(\frac{p+2m+1}{2}\right)\quad\text{for each }a\geq 1.\end{split}$
@@
We fix $a\geq 1$ and denote by $G_{s,p,m,a}(\rho)$ the integrand in the left-hand side of (5.10). Then, $|G_{s,p,m,a}(\rho)|$ is bounded by the function $\sup_{0\leq s\leq 1}|G_{s,p,m,a}(\rho)|$, which is independent of $s$ and bounded on a compact interval $[\delta/a,R]$. Thus, we have
@@
$\begin{split}&\lim_{s\nearrow 1}(1-s)\int_{0}^{R}G_{s,p,m+1,a}(\rho)\,\mathrm{d}\rho\\ &=\lim_{s\nearrow 1}(1-s)\int_{0}^{\delta/a}\rho^{p}\sinh^{2m+2}\rho\left(\frac{-\partial_{\rho}}{\sinh\rho}\right)^{m+1}\mathscr{K}_{\frac{1+sp}{2},a}(\rho)\,\mathrm{d}\rho\\ &\leq\lim_{s\nearrow 1}(1-s)(1+\varepsilon)^{2m+1}\int_{0}^{\delta/a}\rho^{p+2m+1}(-\partial_{\rho})\left(\frac{-\partial_{\rho}}{\sinh\rho}\right)^{m}\mathscr{K}_{\frac{1+sp}{2},a}(\rho)\,\mathrm{d}\rho,\end{split}$

### v3

@@ 4. Extension problem
$P(\rho,y)=C_{4}\,y^{sp}\left(\frac{-\partial_{\rho}}{\sinh\rho}\right)^{\frac{n-1}{2}}\widetilde{K}_{\frac{1+sp}{2},\frac{n-1}{2}}\left(\sqrt{\rho^{2}+y^{2}}\right)$
@@
$P(\rho,y)=C_{4}\,y^{sp}\int_{\rho}^{\infty}\frac{\sinh r}{\sqrt{\pi}\sqrt{\cosh r-\cosh\rho}}\left(\frac{-\partial_{r}}{\sinh r}\right)^{\frac{n}{2}}\widetilde{K}_{\frac{1+sp}{2},\frac{n-1}{2}}\left(\sqrt{r^{2}+y^{2}}\right)\,\mathrm{d}r$
when $n\geq 2$ even; here, $\widetilde{K}_{\nu,a}$ is the function given in (2.4) and
@@
$V(x,y)=\frac{y^{sp}}{2^{sp}\Gamma(\frac{sp}{2})}\int_{0}^{\infty}\int_{\mathbb{H}^{n}}h(t,\rho)f(\xi)\,\mathrm{d}\xi\,e^{-\frac{y^{2}}{4t}}\frac{\mathrm{d}t}{t^{1+\frac{sp}{2}}},$
where $\rho=d(x,\xi)$. Recalling the expression (3.1) for the heat kernel $h(t,\rho)$ and using (2.1), we obtain
@@
$\begin{split}\int_{0}^{\infty}e^{-\frac{(2m-1)^{2}}{4}t-\frac{r^{2}+y^{2}}{4t}}t^{-\frac{5+sp}{2}}\,\mathrm{d}t&=2(2m-1)^{\frac{3+sp}{2}}\widetilde{K}_{\frac{3+sp}{2},\frac{2m-1}{2}}\left(\sqrt{r^{2}+y^{2}}\right)\\ &=4(2m-1)^{\frac{1+sp}{2}}\left(\frac{-\partial_{r}}{r}\right)\widetilde{K}_{\frac{1+sp}{2},\frac{2m-1}{2}}\left(\sqrt{r^{2}+y^{2}}\right).\end{split}$
@@
$\begin{split}V(x,y)&=C_{4}\,y^{sp}\int_{\mathbb{H}^{n}}\left(\frac{-\partial_{\rho}}{\sinh\rho}\right)^{m-1}\int_{\rho}^{\infty}\frac{(-\partial_{r})\widetilde{K}_{\frac{1+sp}{2},\frac{2m-1}{2}}\left(\sqrt{r^{2}+y^{2}}\right)}{\sqrt{\pi}\sqrt{\cosh r-\cosh\rho}}\,\mathrm{d}r\,f(\xi)\,\mathrm{d}\xi\\ &=\int_{\mathbb{H}^{n}}P(d(x,\xi),y)f(\xi)\,\mathrm{d}\xi\end{split}$
@@
Finally, we prove $V(x,0)=f(x)$. Indeed, since the heat kernel $h(t,\rho)$ satisfies
$\int_{\mathbb{H}^{n}}h(t,d(x,\xi))\,\mathrm{d}\xi=1,$
@@
$\begin{split}\int_{\mathbb{H}^{n}}P(d(x,\xi),y)\,\mathrm{d}\xi&=\frac{y^{sp}}{2^{sp}\Gamma(\frac{sp}{2})}\int_{0}^{\infty}\left(\int_{\mathbb{H}^{n}}h(t,d(x,\xi))\,\mathrm{d}\xi\right)e^{-\frac{y^{2}}{4t}}\frac{\mathrm{d}t}{t^{1+\frac{sp}{2}}}\\ &=\frac{y^{sp}}{2^{sp}\Gamma(\frac{sp}{2})}\int_{0}^{\infty}e^{-\frac{y^{2}}{4t}}\frac{\mathrm{d}t}{t^{1+\frac{sp}{2}}}=1.\end{split}$
Now, fix $\varepsilon>0$ and find $\delta>0$ such that $|f(\xi)-f(x)|<\varepsilon$ whenever $d(x,\xi)<\delta$. Then
$|V(x,y)-f(x)|$ $\leq\int_{d(x,\xi)<\delta}P(d(x,\xi),y)|f(\xi)-f(x)|\,\mathrm{d}\xi+2\|f\|_{L^{\infty}}\int_{d(x,\xi)<\delta}P(d(x,\xi),y)\,\mathrm{d}\xi$
$\leq\varepsilon+2\|f\|_{L^{\infty}}\int_{d(x,\xi)<\delta}P(d(x,\xi),y)\,\mathrm{d}\xi.$
Since $P(\rho,y)\to 0$ uniformly for $\rho\geq\delta$ as $y\searrow 0$, we obtain
$\limsup_{y\searrow 0}|V(x,y)-f(x)|\leq\varepsilon.$
Since $\varepsilon>0$ was arbitrary, this shows that $V(x,0)=f(x)$. This concludes that $V$ solves the extension problem (1.5). ∎
@@
$K(\rho)=\left(\frac{-\partial_{\rho}}{\sinh\rho}\right)^{\frac{n-1}{2}}\left(\widetilde{K}_{\frac{1+sp}{2},\frac{n-1}{2}}(\rho)-\widetilde{K}_{\frac{1+sp}{2},\frac{n-1}{2}}\left(\sqrt{\rho^{2}+y^{2}}\right)\right)$
@@
$K(\rho)=\int_{\rho}^{\infty}\frac{\sinh r}{\sqrt{\pi}\sqrt{\cosh r-\cosh\rho}}\left(\frac{-\partial_{r}}{\sinh r}\right)^{\frac{n}{2}}\left(\widetilde{K}_{\frac{1+sp}{2},\frac{n-1}{2}}(r)-\widetilde{K}_{\frac{1+sp}{2},\frac{n-1}{2}}\left(\sqrt{r^{2}+y^{2}}\right)\right)\mathrm{d}r$
@@ 5. Pointwise convergence
as $s\to 1^{-}$, play a key role in the proof of Theorem 1.5. Here, we recall that the constant $c_{n,s,p}$ is given in (1.2).
@@
Let us first consider the case $n=2m+1$ with $m\geq 1$. Since $\lim_{s\nearrow 1}(1-s)|\Gamma(-s)|=1$, we have $c_{n,s,p}C_{1}\leq C(1-s)$ for some $C=C(n,p)>0$. By using Corollary 3.3, we have
(5.3) $\begin{split}0&\leq c_{n,s,p}\int_{R}^{\infty}\mathcal{K}_{n,s,p}(\rho)\sinh^{n-1}\rho\,\mathrm{d}\rho\\ &\lesssim(1-s)\int_{R}^{\infty}\sinh^{2m}\rho\left(\frac{-\partial_{\rho}}{\sinh\rho}\right)^{m}\widetilde{K}_{\frac{1+sp}{2},m}(\rho)\,\mathrm{d}\rho.\end{split}$
@@
(5.4) $\lim_{s\nearrow 1}(1-s)\int_{R}^{\infty}\sinh^{m+a}\rho\left(\frac{-\partial_{\rho}}{\sinh\rho}\right)^{m}\widetilde{K}_{\frac{1+sp}{2},a}(\rho)\,\mathrm{d}\rho=0\quad\text{for each }a>0.$
@@
$\begin{split}\int_{R}^{\infty}\sinh^{1+a}\rho\left(\frac{-\partial_{\rho}}{\sinh\rho}\right)\widetilde{K}_{\frac{1+sp}{2},a}(\rho)\,\mathrm{d}\rho&=a\int_{R}^{\infty}(\sinh^{a}\rho)\rho^{-\frac{1+sp}{2}}K_{\frac{3+sp}{2}}(a\rho)\,\mathrm{d}\rho\\ &\leq a\int_{R}^{\infty}(\sinh^{a}\rho)\rho^{-\frac{1+sp}{2}}K_{\frac{3+p}{2}}(a\rho)\,\mathrm{d}\rho.\end{split}$
@@
$\lim_{s\nearrow 1}(1-s)\int_{R}^{\infty}\sinh^{1+a}\rho\left(\frac{-\partial_{\rho}}{\sinh\rho}\right)\widetilde{K}_{\frac{1+sp}{2},a}(\rho)\,\mathrm{d}\rho=0,$
@@
$\int_{R}^{\infty}\sinh^{m+1+a}\rho\left(\frac{-\partial_{\rho}}{\sinh\rho}\right)^{m+1}\widetilde{K}_{\frac{1+sp}{2},a}(\rho)\,\mathrm{d}\rho$
$=(m+a)\int_{R}^{\infty}\sinh^{m+a-1}\rho\cosh\rho\left(\frac{-\partial_{\rho}}{\sinh\rho}\right)^{m}\widetilde{K}_{\frac{1+sp}{2},a}(\rho)\,\mathrm{d}\rho$
$\quad-\left[\sinh^{m+a}\rho\left(\frac{-\partial_{\rho}}{\sinh\rho}\right)^{m}\widetilde{K}_{\frac{1+sp}{2},a}(\rho)\right]_{R}^{\infty}.$
Note that by Lemma 3.4 and (2.2),
$\sinh^{m+a}\rho\left(\frac{-\partial_{\rho}}{\sinh\rho}\right)^{m}\widetilde{K}_{\frac{1+sp}{2},a}(\rho)\sim\rho^{-\frac{1+sp}{2}}(\sinh\rho)^{a}K_{m+\frac{1+sp}{2}}(a\rho)\sim\rho^{-1-sp/2}$
as $\rho\to\infty$. Thus, taking the limit $\lim_{s\nearrow 1}(1-s)$ yields
$\begin{split}&\lim_{s\nearrow 1}(1-s)\int_{R}^{\infty}\sinh^{m+1+a}\rho\left(\frac{-\partial_{\rho}}{\sinh\rho}\right)^{m+1}\widetilde{K}_{\frac{1+sp}{2},a}(\rho)\,\mathrm{d}\rho\\ &=\lim_{s\nearrow 1}(1-s)(m+a)\int_{R}^{\infty}\sinh^{m+a-1}\rho\cosh\rho\left(\frac{-\partial_{\rho}}{\sinh\rho}\right)^{m}\widetilde{K}_{\frac{1+sp}{2},a}(\rho)\,\mathrm{d}\rho.\end{split}$
@@
$\begin{split}&\lim_{s\nearrow 1}(1-s)\int_{R}^{\infty}\sinh^{m+1+a}\rho\left(\frac{-\partial_{\rho}}{\sinh\rho}\right)^{m+1}\widetilde{K}_{\frac{1+sp}{2},a}(\rho)\,\mathrm{d}\rho\\ &\leq(m+a)(\coth R)\lim_{s\nearrow 1}(1-s)\int_{R}^{\infty}\sinh^{m+a}\rho\left(\frac{-\partial_{\rho}}{\sinh\rho}\right)^{m}\widetilde{K}_{\frac{1+sp}{2},a}(\rho)\,\mathrm{d}\rho=0.\end{split}$
@@
$\begin{split}0&\leq c_{n,s,p}\int_{R}^{\infty}\mathcal{K}_{n,s,p}(\rho)\sinh^{n-1}\rho\,\mathrm{d}\rho\\ &\lesssim(1-s)\int_{R}^{\infty}\sinh^{2m-1}\rho\int_{\rho}^{\infty}\frac{\sinh r}{\sqrt{\cosh r-\cosh\rho}}\left(\frac{-\partial_{r}}{\sinh r}\right)^{m}\widetilde{K}_{\frac{1+sp}{2},\frac{2m-1}{2}}(r)\,\mathrm{d}r\,\mathrm{d}\rho,\end{split}$
@@
(5.7) $\begin{split}\lim_{s\nearrow 1}(1-s)\int_{R}^{\infty}\sinh^{\frac{2m-1}{2}+a}\rho\int_{\rho}^{\infty}\frac{\sinh r}{\sqrt{\cosh r-\cosh\rho}}\left(\frac{-\partial_{r}}{\sinh r}\right)^{m}\widetilde{K}_{\frac{1+sp}{2},a}(r)\,\mathrm{d}r\,\mathrm{d}\rho=0\\ \quad\text{for each }a\geq 1/2.\end{split}$
@@
$\begin{split}&\int_{R}^{\infty}\sinh^{\frac{1}{2}+a}\rho\int_{\rho}^{\infty}\frac{\sinh r}{\sqrt{\cosh r-\cosh\rho}}\left(\frac{-\partial_{r}}{\sinh r}\right)\widetilde{K}_{\frac{1+sp}{2},a}(r)\,\mathrm{d}r\,\mathrm{d}\rho\\ &\leq a\int_{R}^{M/a}\sinh^{\frac{1}{2}+a}\rho\int_{\rho}^{\infty}\frac{1}{\sqrt{\cosh r-\cosh\rho}}r^{-\frac{1+sp}{2}}K_{\frac{3+p}{2}}(ar)\,\mathrm{d}r\,\mathrm{d}\rho\\ &\quad+a\int_{M/a}^{\infty}\sinh^{\frac{1}{2}+a}\rho\int_{\rho}^{\infty}\frac{1}{\sqrt{\cosh r-\cosh\rho}}r^{-\frac{1+sp}{2}}K_{\frac{3+p}{2}}(ar)\,\mathrm{d}r\,\mathrm{d}\rho=:J_{1}+J_{2}.\end{split}$
@@
$\begin{split}&\int_{R}^{\infty}\sinh^{\frac{2m+1}{2}+a}\rho\int_{\rho}^{\infty}\frac{\sinh r}{\sqrt{\cosh r-\cosh\rho}}\left(\frac{-\partial_{r}}{\sinh r}\right)^{m+1}\widetilde{K}_{\frac{1+sp}{2},a}(r)\,\mathrm{d}r\,\mathrm{d}\rho\\ &=\int_{R}^{\infty}\sinh^{\frac{2m+1}{2}+a}\rho\left(\frac{-\partial_{\rho}}{\sinh\rho}\right)\int_{\rho}^{\infty}\frac{\sinh r}{\sqrt{\cosh r-\cosh\rho}}\left(\frac{-\partial_{r}}{\sinh r}\right)^{m}\widetilde{K}_{\frac{1+sp}{2},a}(r)\,\mathrm{d}r\,\mathrm{d}\rho.\end{split}$
@@
$\begin{split}&\lim_{s\nearrow 1}(1-s)\int_{R}^{\infty}\sinh^{\frac{2m+1}{2}+a}\rho\int_{\rho}^{\infty}\frac{\sinh r}{\sqrt{\cosh r-\cosh\rho}}\left(\frac{-\partial_{r}}{\sinh r}\right)^{m+1}\widetilde{K}_{\frac{1+sp}{2},a}(r)\,\mathrm{d}r\,\mathrm{d}\rho\\ &\leq C\lim_{s\nearrow 1}(1-s)\int_{R}^{\infty}\sinh^{\frac{2m-1}{2}+a}\rho\int_{\rho}^{\infty}\frac{\sinh r}{\sqrt{\cosh r-\cosh\rho}}\left(\frac{-\partial_{r}}{\sinh r}\right)^{m}\widetilde{K}_{\frac{1+sp}{2},a}(r)\,\mathrm{d}r\,\mathrm{d}\rho\end{split}$
@@
(5.9) $\begin{split}&\lim_{s\nearrow 1}(1-s)\int_{0}^{R}\rho^{p}\sinh^{2m}\rho\left(\frac{-\partial_{\rho}}{\sinh\rho}\right)^{m}\widetilde{K}_{\frac{1+sp}{2},m}(\rho)\,\mathrm{d}\rho\\ &=\frac{2^{m-1}}{p}\left(\frac{2}{m}\right)^{\frac{p+1}{2}}\Gamma\left(\frac{p+2m+1}{2}\right)\end{split}$
@@
(5.10) $\begin{split}\lim_{s\nearrow 1}(1-s)\int_{0}^{R}\rho^{p}\sinh^{2m}\rho\left(\frac{-\partial_{\rho}}{\sinh\rho}\right)^{m}\widetilde{K}_{\frac{1+sp}{2},a}(\rho)\,\mathrm{d}\rho\\ =\frac{2^{m-1}}{p}\left(\frac{2}{a}\right)^{\frac{p+1}{2}}\Gamma\left(\frac{p+2m+1}{2}\right)\quad\text{for each }a\geq 1.\end{split}$
@@
We fix $a\geq 1$ and denote by $G_{s,p,m,a}(\rho)$ the integrand in the left-hand side of (5.10). Then, $|G_{s,p,m,a}(\rho)|$ is bounded by the function $\sup_{0\leq s\leq 1}|G_{s,p,m,a}(\rho)|$, which is independent of $s$ and bounded on a compact interval $\delta/a,R]$. Thus, we have
@@
$\begin{split}&\lim_{s\nearrow 1}(1-s)\int_{0}^{R}G_{s,p,m+1,a}(\rho)\,\mathrm{d}\rho\\ &=\lim_{s\nearrow 1}(1-s)\int_{0}^{\delta/a}\rho^{p}\sinh^{2m+2}\rho\left(\frac{-\partial_{\rho}}{\sinh\rho}\right)^{m+1}\widetilde{K}_{\frac{1+sp}{2},a}(\rho)\,\mathrm{d}\rho\\ &\leq\lim_{s\nearrow 1}(1-s)(1+\varepsilon)^{2m+1}\int_{0}^{\delta/a}\rho^{p+2m+1}(-\partial_{\rho})\left(\frac{-\partial_{\rho}}{\sinh\rho}\right)^{m}\widetilde{K}_{\frac{1+sp}{2},a}(\rho)\,\mathrm{d}\rho,\end{split}$